OpenStudy (anonymous):

Graph the line with a slope of 1/3 that contains the point (−3, 5).

OpenStudy (anonymous):

y = mx+b y= 5 x= -3 m= 1/3 so: 5 = (1/3) (-3) + b equation 5 = -3/3 + b simplified add (3/3) to both sides

OpenStudy (anonymous):

And you can do the rest.

OpenStudy (anonymous):

which is adding the 3/3 to 5

OpenStudy (anonymous):

k thanks